OpenStudy (anonymous):

6*e^x = 25 solve for x

OpenStudy (anonymous):

take ln on both sides: ln6*e^x = ln25

OpenStudy (anonymous):

x= ln*25/6. first divide bothsides by 6 then take the ln of both sides to get rid of the e

OpenStudy (anonymous):

Use Ln rules to get: ln6 + ln(e^x) = ln25 ln6 + x*ln(e) = ln25

OpenStudy (anonymous):

ln6 + x = ln25 x = ln25-ln6 which simplifies to: x = ln(25/6) I confirm anliker's answer.